# Exploring the Alignment Landscape: LLMs and Geometric Deep Models in Protein Representation

### v1

@@ 3.4.1 GDM Dimension Analysis (RQ4): Does increasing the GDM dimension improve alignment performance?
After retraining the GearNet model, we conducted the experiment outlined in Research Question 1. This involved training the projection head for each side of the model pair and evaluating the performance using our defined metric. As illustrated in Figure 3b, there is a clear trend indicating that as the dimensionality of the GearNet’s embeddings increases, the alignment score of the model pairs improves. This pattern is consistent regardless of which LLM is used. Additionally, we observe that model pairs with larger LLMs, such as LLaMa3.1-70B, tend to achieve higher alignment scores. An interesting observation is that when the GearNet embedding dimension is 256 or smaller, model pairs with Gemma2-2B and LLaMa3.1-8B perform comparably, with each occasionally outperforming the other. However, once the GearNet dimension exceeds 512, model pairs with LLaMa3.1-8B consistently outperform those with Gemma2-2B. This result is consistent with our findings in Research Question 1, where we observed that as the LLM’s parameters increased, its representation captured richer information, leading to better alignment. Similarly, in this experiment, the larger GearNet dimensions introduce richer, more detailed representations of the protein graph, resulting in enhanced alignment scores. This suggests that increasing the complexity of the GDM, like increasing the parameters in the LLM, helps the model capture more intricate biological information.
Figure 5: Number of Projection Head Layer Analysis. The figure consists of three 3D graphs: model pairs with GearNet are shown on the left, model pairs with GVP in the middle, and model pairs with ScanNet on the right. Each graph has three axes: the x-axis (horizontal axis in the foreground) represents the LLM used in the model pair, the y-axis (horizontal axis on the left) represents the number of layers in the projection head, and the z-axis (vertical axis) represents the alignment score.
3.4.2 Projection Head Analysis (RQ5): Does adding layers to the GDM’s projection head enhance alignment performance?
In previous experiments, all projection heads consisted of a single linear layer. In this section, we explore whether increasing the complexity of the GDM’s projection head by adding more layers improve alignment performance. To investigate this, we added one and two additional linear layers to the GDM’s projection head while keeping the LLM’s projection head structure unchanged. This results in GDM projection heads with 2 or 3 linear layers, with ReLU activation functions between layers. We then re-trained the projection heads and evaluated alignment performance as in previous experiments. Further details of the implementation are provided in Appendix A.7.
As shown in Figure 5, the x-axis represents different model pairs, the y-axis represents the number of layers, and the z-axis represents the alignment score. The results indicate that increasing the number of layers in the projection head consistently improves alignment scores for the majority of model pairs. Notably, projection heads with 2 layers significantly outperform the 1-layer architecture across all model pairs. However, adding a third layer provides only marginal improvements and, in most cases, actually decreases the alignment score, especially for model pairs involving ScanNet. The improvement from 1 to 2 layers is substantial, suggesting that a 2-layer architecture provides a significant boost in alignment capability. However, the minimal gains observed when moving from 2 to 3 layers indicate diminishing returns with increased complexity. Thus, the 2-layer architecture appears to strike the best balance between model complexity and performance improvement. This suggests that the number of projection head layers should be carefully determined based on empirical validation for each specific model combination, rather than assuming a one-size-fits-all approach. Figure 8b shows the continuous graph analysis of model pairs with GAT due to the page limit.
3.4.3 LLM Fine-Tuning Analysis (RQ6): Does fine-tuning an LLM on protein data enhance alignment performance?
In previous experiments, all model pairs used pre-trained LLMs directly downloaded from Huggingface. In this section, we fine-tuned all LLMs on the dataset constructed during the FASTA preprocessing stage, where we generated a detailed text description containing all relevant information for each protein. The fine-tuning process involved framing the task as a question-answering problem, where the input was a prompt such as, “What is the protein: XXX?” and the expected output was the corresponding text description. Further implementation details are provided in Appendix A.8. After fine-tuning the LLMs, we trained the 1-layer projection heads for the each model pairs and evaluated the alignment performance as before. As shown in Figure 6, the results clearly show that fine-tuning improves alignment scores across all model pairs. This indicates that providing the LLM with domain-specific knowledge enhances its ability to align with GDM representations. These findings suggest that LLMs fine-tuned on domain-specific data are more capable of generating representations that align effectively with other modalities. This underscores the importance of domain-specific LLMs when developing MMLMs, as they facilitate better alignment with non-text modalities such as graphs.
@@
Figure 6: LLM Fine-tuned on Protein Data Analysis. The figure is divided into three sections: model pairs with Gemma2-2B on the left, model pairs with LLaMa3.1-8B in the middle, and model pairs with LLaMa3.1-70B on the right. The y-axis represents the alignment score, and the x-axis indicates whether the LLM is fine-tuned or not. Each line represents the trend in alignment scores for a specific model pair, from the original (pre-fine-tuning) to post-fine-tuning.
4 Discussions
@@
4.1 Alignment Observations and Takeaways
@@
4.2 Suggestions for Developing Multimodal LLMs
@@
5 Conclusions
@@
In this paper, we conducted a comprehensive analysis of multimodal alignment between LLMs and GDMs in the protein domain, addressing six key research questions. Our study analyzed the models alignment and proteins alignment, such as the types of models that align well together, and specific factors that influence alignment performance, such as protein rarity and model complexity. Through these experiments, we uncovered significant challenges in multimodal alignment and provided valuable insights into its underlying mechanics. Our findings offer practical guidelines for designing more effective multimodal models. We highlighted the importance of accounting for nuanced and homologous relationships between proteins, emphasizing that simple mappings may overlook critical similarities. We demonstrated the importance of leveraging 3D geometric information in GDMs, maximizing embedding dimensions in LLMs, and fine-tuning LLMs with domain-specific knowledge. Additionally, we identified optimal layers for projection heads to enhance alignment without unnecessary complexity. These insights pave the way for future research and development, helping to create more robust and accurate multimodal systems that can better understand and represent protein data.
Acknowledgment

### v2

@@ Results
After retraining the GearNet model, we conducted the experiment outlined in Research Question 1. This involved training the projection head for each side of the model pair and evaluating the performance using our defined metric. As illustrated in Figure 2C, there is a clear trend indicating that as the dimensionality of the GearNet’s embeddings increases, the alignment score of the model pairs improves. This pattern is consistent regardless of which LLM is used. Additionally, we observe that model pairs with larger LLMs, such as LLaMa3.1-70B, tend to achieve higher alignment scores. An interesting observation is that when the GearNet embedding dimension is 256 or smaller, model pairs with Gemma2-2B and LLaMa3.1-8B perform comparably, with each occasionally outperforming the other. However, once the GearNet dimension exceeds 512, model pairs with LLaMa3.1-8B consistently outperform those with Gemma2-2B. This result is consistent with our findings in Research Question 1, where we observed that as the LLM’s parameters increased, its representation captured richer information, leading to better alignment. Similarly, in this experiment, the larger GearNet dimensions introduce richer, more detailed representations of the protein graph, resulting in enhanced alignment scores. This suggests that increasing the complexity of the GDM, like increasing the parameters in the LLM, helps the model capture more intricate biological information.
@@
Figure 4: Strategies for Improvement. (A) The figure consists of three 3D graphs: model pairs with GearNet are shown on the left, model pairs with GVP in the middle, and model pairs with ScanNet on the right. Each graph has three axes: the x-axis (horizontal axis in the foreground) represents the LLM used in the model pair, the y-axis (horizontal axis on the left) represents the number of layers in the projection head, and the z-axis (vertical axis) represents the alignment score. (B) The figure is divided into three sections: Gemma2-2B on the left, LLaMa3.1-8B in the middle, and LLaMa3.1-70B on the right. The y-axis represents the alignment score, and the x-axis indicates whether the LLM is fine-tuned or not. Each line represents the trend in alignment scores for a specific model pair, from the original (pre-fine-tuning) to post-fine-tuning.
Projection Head Analysis (RQ5): Does adding layers to the GDM’s projection head enhance alignment performance?
In previous experiments, all projection heads consisted of a single linear layer. In this section, we explore whether increasing the complexity of the GDM’s projection head by adding more layers improve alignment performance. To investigate this, we added one and two additional linear layers to the GDM’s projection head while keeping the LLM’s projection head structure unchanged. This results in GDM projection heads with 2 or 3 linear layers, with ReLU activation functions between layers. We then re-trained the projection heads and evaluated alignment performance as in previous experiments. Further details of the implementation are provided in Supplemental Note 7.
As shown in Figure 4A, the x-axis represents different model pairs, the y-axis represents the number of layers, and the z-axis represents the alignment score. The results indicate that increasing the number of layers in the projection head consistently improves alignment scores for the majority of model pairs. Notably, projection heads with 2 layers significantly outperform the 1-layer architecture across all model pairs. However, adding a third layer provides only marginal improvements and, in most cases, actually decreases the alignment score, especially for model pairs involving ScanNet. The improvement from 1 to 2 layers is substantial, suggesting that a 2-layer architecture provides a significant boost in alignment capability. However, the minimal gains observed when moving from 2 to 3 layers indicate diminishing returns with increased complexity. Thus, the 2-layer architecture appears to strike the best balance between model complexity and performance improvement. This suggests that the number of projection head layers should be carefully determined based on empirical validation for each specific model combination, rather than assuming a one-size-fits-all approach. Figure S6B shows the continuous graph analysis of model pairs with GAT due to the page limit.
LLM Fine-Tuning Analysis (RQ6): Does fine-tuning an LLM on protein data enhance alignment performance?
In previous experiments, all model pairs used pre-trained LLMs directly downloaded from Huggingface. In this section, we fine-tuned all LLMs on the dataset constructed during the FASTA preprocessing stage, where we generated a detailed text description containing all relevant information for each protein. The fine-tuning process involved framing the task as a question-answering problem, where the input was a prompt such as, “What is the protein: XXX?” and the expected output was the corresponding text description. Further implementation details are provided in Supplemental Note 8. After fine-tuning the LLMs, we trained the 1-layer projection heads for the each model pairs and evaluated the alignment performance as before. As shown in Figure 4B, the results clearly show that fine-tuning improves alignment scores across all model pairs. This indicates that providing the LLM with domain-specific knowledge enhances its ability to align with GDM representations. These findings suggest that LLMs fine-tuned on domain-specific data are more capable of generating representations that align effectively with other modalities. This underscores the importance of domain-specific LLMs when developing multimodal models, as they facilitate better alignment with non-text modalities such as graphs.
Discussion
@@
Alignment Observations and Takeaways
@@
Protein-focused MLLMs Design Recommendations
@@
Table 2: Protein-focused MLLMs Performance on Protein Description Task
Models Gemma2-2B LLaMa3.1-8B LLaMa3.1-70B
ROUGE BLEU ROUGE BLEU ROUGE BLEU
GearNet 0.3761 0.1914 0.4164 0.2125 0.4849 0.2574
GAT 0.1256 0.0903 0.1278 0.0916 0.1315 0.1009
GVP 0.1364 0.0952 0.1420 0.0978 0.1536 0.1058
ScanNet 0.4203 0.2057 0.4614 0.2392 0.5127 0.2620
Protein-focused MLLMs Performance and Applications
We have explored the alignment landscape between LLMs and GDMs, proposing several strategies to enhance their alignment. However, a critical question arises: ‘How does this alignment contribute to protein modeling or real-world applications?’ To address this, we developed protein-focused MLLMs by aligning each LLM-GDM model pair and further fine-tuning the projection head within the protein-focused MLLM. Existing research has demonstrated that the degree of alignment in multimodal models is strongly correlated with their performance on specific downstream tasks. Additionally, alignment levels play a significant role in mitigating hallucinations, where models generate plausible but factually incorrect information 26, 27. To evaluate the impact of alignment on both performance and hallucination, we selected the protein description task and employed ROUGE 28 and BLEU 29 metrics. Detailed methodology for building protein-focused MLLMs and example outputs are provided in Supplemental Note 9.
As shown in Table 2, protein-focused MLLMs utilizing GAT and GVP as their GDMs exhibit poor performance on the protein description task. In contrast, protein-focused MLLMs that employ GearNet and ScanNet as their GDMs achieve superior results. Additionally, we observe a significant improvement in performance as the underlying LLM becomes more powerful (e.g., from Gemma2-2B to LLaMa3.1-70B). This finding aligns with our results from Research Question 1. Furthermore, ROUGE scores are generally higher than BLEU scores. This discrepancy arises because BLEU emphasizes precision—measuring how much of the generated output matches the ground truth—whereas ROUGE focuses on recall, which captures how much of the ground truth is reflected in the generated output. Upon examining the outputs generated by protein-focused MLLMs (see Figure S7b), we discovered that models often generate additional protein-related knowledge not included in our ground truth. These extra details, while potentially accurate, are not reflected in the ground truth, resulting in lower BLEU scores compared to ROUGE. This highlights the need for a more comprehensive protein description benchmark that encompasses a broader and more diverse range of protein knowledge than ours, paving the way for more robust evaluations in future research.
To further illustrate the relationship between alignment levels and hallucination, we present examples of protein-focused MLLMs generated descriptions in Figure S7b for a popular protein ID 4NWH. When the model’s alignment is low, most of the generated content consists of hallucinations. This occurs because the model fails to align the protein representation with the actual protein, instead aligning it with unrelated concepts. For example, models using GAT and GVP as GDMs generate hallucinated and irrelevant content, such as ‘The protein is a protein…’ or ‘It is a large molecule made up of long chains of amino acids.’ Although these statements are technically correct, they lack meaningful insights and are as unhelpful as hallucinated content. In contrast, as alignment improves, the level of hallucination decreases significantly. This improvement is evident when comparing models that employ GearNet or ScanNet as GDMs and LLaMa3.1-70B as the LLM to those with lower alignment levels. These findings validate our experimental results and corroborate existing research, demonstrating a strong correlation between higher alignment, better downstream task performance, and reduced hallucination.
Table 3: Model Pairs Alignment Performance (Reweight)
Models Trained
Gemma2-2B LLaMa3.1-8B LLaMa3.1-70B
GearNet 0.2958+0.0260 0.3211+0.0425 0.3515+0.0491
GAT -0.0533+0.0048 -0.0700+0.0282 -0.0610+0.0164
GVP 0.0774+0.0192 0.0965+0.0051 0.1408+0.0120
ScanNet 0.3223+0.0367 0.3375+0.0478 0.3856+0.0496
Rare Protein Alignment Strategies
In this section, we address the limitations and challenges posed by current protein datasets, particularly the imbalance in the study of popular versus rare proteins. To tackle this rare protein challenge, we propose solutions from two perspectives: improving model-pair alignment and enhancing the downstream task performance of protein-focused MLLMs.
To improve model-pair alignment, we employ a reweighting technique during the training of the projection head. This technique prioritizes rare proteins by assigning higher penalties to errors associated with them. Detailed information on this method can be found in Supplemental Note 10. Beyond alignment, we introduce a retrieval-augmented approach to enhance protein-focused MLLMs’ performance on the protein description task. Using the reweighted projection head, we follow the methodology described in Supplemental Note 9 to construct the protein-focused MLLM. During input stage, our method retrieves the top $k$ most similar proteins to the input protein and augments the original input with this retrieved information. Further details on the retrieval method are provided in Supplemental Note 11.
Table LABEL:table:alignment_reweight presents the alignment scores of model pairs after applying the reweighting technique. Original scores from Table LABEL:table:alignment_comparison are shown in black, while improvements are in italic. These improvements demonstrate the effectiveness of reweighting in addressing the imbalance between rare and popular proteins. While our classification of rare and popular proteins was not completely, our results suggest that solving the rare protein problem is a promising direction for improving model-pair alignment. Future research should focus on creating a more comprehensive dataset that explicitly distinguishes between rare and popular proteins, enabling further advancements in alignment strategies.
Table 4 compares the downstream performance of protein-focused MLLMs before and after applying the retrieval method. Both the reweighting technique alone and the combination of reweighting with retrieval show practical improvements. However, the gains from reweighting alone are limited, as the technique only targets a subset of rare proteins during training. Consequently, a significant portion of rare proteins is penalized similarly to popular proteins. This finding underscores the need for a more detailed classification of rare and popular proteins to optimize reweighting strategies. In contrast, the retrieval method demonstrates significant performance gains for protein-focused MLLMs, particularly with $k=3$ and $k=5$. Figure S10 illustrates example outputs for rare protein 3I1A before and after applying retrieval with $k=3$. While reweighting improves overall model-pair alignment, it does not fully mitigate hallucinations for rare protein inputs. However, after applying the retrieval method, hallucination rates are substantially reduced, even when compared to the outputs for popular proteins shown in Figure S7b. Moreover, the format and structure of the generated descriptions align more closely with our ground truths, explaining the observed improvements in ROUGE and BLEU scores. These results highlight the importance of a comprehensive ground truth for protein descriptions, as model performance, retrieval augmentation, and evaluation metrics all heavily depend on its quality.
On the other hand, increasing $k$ to 10 results in decreased model performance. We attribute this behavior to the retrieval algorithm retrieving less similar proteins as $k$ increases. These less similar proteins introduce noise or conflicting information, which negatively impacts model performance.
Table 4: Protein-focused MLLMs Performance Before and After Retrieval
Gemma2-2B LLaMa3.1-8B LLaMa3.1-70B
model ROUGE BLEU ROUGE BLEU ROUGE BLEU
$reweight$ GearNet 0.3761+0.0108 0.1914+0.0049 0.4164+0.0141 0.2125+0.0087 0.4849+0.0255 0.2574+0.0074
GAT 0.1256+0.0006 0.0903+0.0002 0.1278+0.0014 0.0916+0.0001 0.1315+0.0012 0.1009+0.0008
GVP 0.1364+0.0079 0.0952+0.0014 0.1420+0.0031 0.0978+0.0005 0.1536+0.0046 0.1058+0.0010
ScanNet 0.4203+0.0133 0.2057+0.0053 0.4614+0.0174 0.2392+0.0067 0.5127+0.0247 0.2620+0.0081
$k=3$ GearNet 0.3761+0.1439 0.1914+0.3167 0.4164+0.1343 0.2125+0.3128 0.4849+0.0985 0.2574+0.3207
GAT 0.1256+0.0044 0.0903+0.0131 0.1278+0.0061 0.0916+0.0227 0.1315+0.0047 0.1009+0.0141
GVP 0.1364+0.0408 0.0952+0.0819 0.1420+0.0381 0.0978+0.0742 0.1536+0.0360 0.1058+0.0654
ScanNet 0.4203+0.1231 0.2057+0.3006 0.4614+0.1138 0.2392+0.3341 0.5127+0.1280 0.2620+0.3565
$k=5$ GearNet 0.3761+0.1697 0.1914+0.3418 0.4164+0.1557 0.2125+0.3374 0.4849+0.1183 0.2574+0.3443
GAT 0.1256+0.0071 0.0903+0.0147 0.1278+0.0068 0.0916+0.0253 0.1315+0.0052 0.1009+0.0169
GVP 0.1364+0.0536 0.0952+0.0861 0.1420+0.0579 0.0978+0.0829 0.1536+0.0415 0.1058+0.0834
ScanNet 0.4203+0.1678 0.2057+0.3519 0.4614+0.1492 0.2392+0.3506 0.5127+0.1497 0.2620+0.3904
$k=10$ GearNet 0.3761+0.1127 0.1914+0.2491 0.4164+0.1136 0.2125+0.2382 0.4849+0.0902 0.2574+0.2965
GAT 0.1256+0.0036 0.0903+0.0124 0.1278+0.0054 0.0916+0.0188 0.1315+0.0031 0.1009+0.0136
GVP 0.1364+0.0333 0.0952+0.0616 0.1420+0.0483 0.0978+0.0709 0.1536+0.0375 0.1058+0.0713
ScanNet 0.4203+0.0910 0.2057+0.2743 0.4614+0.1028 0.2392+0.3180 0.5127+0.0844 0.2620+0.3521
Conclusions
@@
In this paper, we conducted a comprehensive analysis of multimodal alignment between LLMs and GDMs in the protein domain, addressing six key research questions. Our study analyzed the models alignment and proteins alignment, such as the types of models that align well together, and specific factors that influence alignment performance, such as protein rarity and model complexity. Through these experiments, we uncovered significant challenges in multimodal alignment and provided valuable insights into its underlying mechanics. Our findings offer practical guidelines for designing more effective multimodal models. We highlighted the importance of accounting for nuanced and homologous relationships between proteins, emphasizing that simple mappings may overlook critical similarities. We demonstrated the importance of leveraging 3D geometric information in GDMs, maximizing embedding dimensions in LLMs, and fine-tuning LLMs with domain-specific knowledge. Additionally, we identified optimal layers for projection heads to enhance alignment without unnecessary complexity. Furthermore, we translated the alignment between LLMs and GDMs into practical applications by developing protein-focused MLLMs and evaluating them on the protein description task. The downstream performance of our models confirms that higher alignment leads to better performance and reduced hallucinations. Through our experiments, we also highlighted the need for a more comprehensive protein dataset that categorizes proteins based on their popularity and rarity, while also incorporating detailed and diverse descriptions on each protein. These insights pave the way for future research and development, helping to create more robust and accurate multimodal systems that can better understand and represent protein data.
Methods
Models: In this paper, we selected four state-of-the-art GDMs specialized in the protein domain, as well as three pre-trained LLMs. Each model has a distinct embedding dimensionality, adding diversity to our experiments. For GDMs, we used the following models: GAT with an embedding size of 64 12, ScanNet with an embedding size of 128 30, GVP with an embedding size of 148 16, and GearNet with an embedding size of 3072 31. All GDMs were pre-trained and loaded directly from their respective official repositories to ensure consistent and stable performance. For LLMs, we selected Gemma2-2B with an embedding size of 2304 32, LLaMa3.1-8B with an embedding size of 4096 33, and LLaMa3.1-70B with an embedding size of 8192. All LLMs were loaded directly from Huggingface’s repository.
Dataset: We downloaded protein data from the RCSB PDB database 34. For this study, we randomly selected 20,000 protein samples. Each protein has two modalities: a text-based description from the FASTA file and a 3D structural graph modality from the PDB file. We randomly split the dataset into 80% for training, 10% for validation, and 10% for testing. A more detailed statistic of the dataset is shown in Supplemental Note 1.
Step I: Data Preprocessing
FASTA Preprocessing
Since the selected LLMs are pre-trained on general text data without specific knowledge of proteins, we convert the FASTA file into a clear and detailed text description that the LLMs can understand. This is shown in Figure 1 Data Preprocessing upper part. The process requires stripping unnecessary information and retaining only the essential content. A typical FASTA file contains several key details: protein ID, chains, molecule name, organism, and the amino acid sequence. While we consider most of this information important, the amino acid sequence itself is often lengthy and may introduce noise, so we exclude it from the text description and instead include the sequence length. Additionally, the FASTA file allows us to determine whether a protein consists of a single chain or multiple chains. For single-chain proteins, we use regular expressions to extract relevant details. However, for multi-chain proteins, the complexity increases, and regular expressions alone often fail to capture all information. In such cases, we employ GPT-4o to generate accurate descriptions, extracting the chains and corresponding organisms. The GPT-4o prompt is provided in Supplemental Note 2.
After processing, each protein is described with text similar to: “The protein structure [protein_id] has a sequence length of [number] amino acids. Here is more information: The protein structure [protein_id] involves the following chains: [chains]. The protein is named [protein_name] and is derived from the organism [organism].” We manually validated the GPT-generated descriptions for all multi-chain proteins and found them to be accurate compared to the original FASTA files. Examples of protein descriptions are shown in Supplemental Note 3.
PDB Preprocessing
[{ ID: ID,
seq: TQDCSFQHSP...’,
coords: [[[74.46, 58.25, -21.65],...],...]}...]
Step II: Latent Representation Extraction
Once the data has been processed, the next step is to feed it into the respective models for latent representation extraction. This process is shown in Figure 1 “Representation Extraction” part.
LLMs Representation
GDMs Representation
  • GearNet: GearNet encodes protein structure information through a relational graph neural network (GNN) and an edge message-passing mechanism. The input to the model consists of a residue-level relational graph constructed from the 3D structure of the protein, which Gearnet’s code will derived from the input PDB file automatically. The detail is shown in Supplemental Note 4.1. In the end, to obtain the final protein representation, the features from all hidden layers are concatenated, which results in the final protein representation has the shape $[1,3072]$.
[…]
  • GAT: Previously, we already processed the protein PDB file into the format that GAT supports. Therefore, graph $G$ is fed into a GAT model to compute latent representations. The detail is shown in Supplemental Note 4.4. In GAT architecture, they use 8 attention heads, with each head computing 8 features, leading to a concatenated output of size $64$ for each node, which results in protein’s final representation size of [1, node_size, 64]. Then we simply perform average pooling to remove the ‘node_size’ dimension. This process yields the latent representation of the protein with size of [1, 64], which will be stored and used later.
Step III: Representation Alignment
Contrastive Loss Function
To maximize the cosine similarity of positive pairs to 1 and minimize that of negative pairs to 0, we apply a contrastive loss function during training. Specifically, we use a modified version of the InfoNCE loss 35. Let $g_{i}$ be the projected graph representation for protein $i$, $t_{j}$ be the projected text representation for protein $j$, and $\text{sim}(g_{i},t_{j})$ represent the cosine similarity between $g_{i}$ and $t_{j}$, adjusted to the range $[0,1]$ by using $\frac{\text{sim}(g_{i},t_{j})+1}{2}$. The temperature parameter $\tau$ is set to 0.2 36, and the batch size $B$ is 32. For each protein $i$, the positive pair similarity is given by $\text{sim}^{+}_{i}=\frac{\text{sim}(g_{i},t_{i})+1}{2}$, and the negative pair similarities are $\text{sim}^{-}_{ij}=\frac{\text{sim}(g_{i},t_{j})+1}{2}$ for all $j\neq i$ within the batch. After calculating these similarities, we apply exponentiation and scale them by the temperature $\tau$:
This overall loss function averages the loss over all proteins in the batch, effectively encouraging high similarity for positive pairs and low similarity for negative pairs. The detail of reweighting technique is shown in Supplemental Note 10.
Implementation Details
Resource Availability
Lead Contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Dong Shu (dongshu2024@u.northwestern.edu).
Materials Availability
This study did not generate new materials.
Data and Code Availability
  • Our data was downloaded directly from the RCSB PDB database https://www.rcsb.org/ and are publicly available. All other data reported in this paper will be shared by the lead contact upon request.
  • The code to reproduce the results of this article can be found at https://github.com/Tizzzzy/LLM-GDM-alignment and is publicly available. Additionally, we have archived the code on Zenodo 37, where it can be accessed via DOI: 10.5281/zenodo.14934851.
  • Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
Acknowledgments